OpenStudy (sushi121212):

Find the absolute value of the complex number: -7-i√11

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

the absolute value is the length of the hypotenuse

OpenStudy (sushi121212):

So I have to do the Pythagorean theorem?

OpenStudy (unklerhaukus):

yes!

OpenStudy (unklerhaukus):

A^2 = (-7)^2 + (-√11)^2 A = √( (-7)^2 + (-√11)^2 )

OpenStudy (sushi121212):

So it would be: A = √(49+11) A = 2√15 Or is it subtracting?

OpenStudy (unklerhaukus):

that's right!

OpenStudy (sushi121212):

Oh!! Thank you so much!!!

OpenStudy (unklerhaukus):

the squares mean that the negative signs all go away